# To reset, or not to reset – that is the question

### v1

@@ IV Recovering time-like distance by squeezing two QEC rounds into one
In this section, we show that QEC performance can be improved without using unconditional resets by effectively squeezing two QEC rounds into one. This method, however, requires $\approx 50\%$ additional qubits. We appoint two auxiliary qubits for each stabiliser, as shown in Figure 11(a). Note that the auxiliary qubits of the $X$-/$Z$-type stabilisers are arranged horizontally/vertically so that we avoid so-called “bad hook” errors. Therefore, the hardware needs the Cairo pentagonal connectivity [33]. The main idea is that while we measure the auxiliary qubits of one stabiliser type, we can execute unitary gates on the qubits of the other type and also have plenty of time to obtain two independent measurement outcomes for each stabiliser. Figure 11(b)–(c) illustrates this. For instance, the unitary parts of this circuit for a $Z$-type stabiliser are as follows. First, apply two layers of CZ gates between the top auxiliary and the top two data qubits and at the same time similarly for the bottom three qubits. Then, swap the auxiliary qubits, and finally repeat the two layers of CZ gates. In this way, we entangle the two auxiliary qubits with the data qubits independently and so measuring each provides independent stabiliser measurements. Since executing these unitary gates takes only $400$ ns in total (when compiled to the gates of Table 1), one QEC round with this circuit takes $1000$ ns. This is only $160$ ns longer than for the standard no-reset syndrome extraction; however, we obtain each stabiliser outcome twice.
@@
We simulated these circuits in order to compare the performance to the previously discussed syndrome extraction circuits. We applied the same principles for the number of shots as in Sections II and III. We present the quantum memory results in Figure 13 and see that again the performance is slightly diminished from the no-reset case, though not as much as with the error-spreading circuits. We attribute this to the increased number of two-qubit gates used in these syndrome extraction circuits. However, as we see from the stability time overhead plots in Figure 14(b), we need fewer QEC rounds for the round-squeezing circuit to avoid undetectable time-like logical failures and so we need fewer QEC rounds for quantum memory as well, making the gap in Figure 13 smaller. For instance, with $p=0.001$ we have $R_{\mathrm{sqz}}\approx 0.65$. Since execution times per QEC round differ, we need only $\approx 55\%$ of the number of QEC rounds with the round-squeezing circuit – slightly more than half as many as for the standard no-reset case. Based on this, we can conclude that this circuit improves on the QEC performance of both the standard and error-spreading no-reset circuits, though with the use of additional physical qubits.
From comparing Figure 14(b) and Figure 5(c), we can further conclude that this circuit is a good alternative to implementing unconditional resets, even with instantaneous reset, at least for the range of physical error parameters we consider. We expect that, with much lower physical error probability, this would no longer be the case.
@@
Figure 14: Comparison of stability experiment performances in the no-reset scheme using standard (triangles and dashed lines) and round-squeezing (diamonds and dotted lines) circuits. Colours indicate different physical error probabilities: $p=10^{-2}$ (blue), $p=10^{-2.5}$ (orange), $p=10^{-3}$ (purple). (a) Comparison of logical failure probabilities for a $4\times 4$ stability patch. Error bars as in Figure 5(a) are also shown as shaded areas. Note that the standard $p=10^{-3}$ and round-squeezing $p=10^{-2.5}$ plots overlap. (b) The time overhead $R_{\mathrm{sqz}}$ that the round-squeezing circuit requires when compared to the no-reset standard circuit.
@@ V Conclusion
Nevertheless, with current best-in-class devices operating close to the QEC threshold and focused more on memory experiments, our simulations indicated that not resetting qubits can result in better performance. This justifies the approach taken in recent proof-of-principle experiments [6, 8] and offers encouragement for near-term experiments to continue this practice. For those keen to avoid developing unconditional-reset technology, we presented two alternative syndrome extraction circuits that provide additional protection against measurement errors. These alternative circuits are interesting near-term experiments and potential long-term solutions. Whether our alternative circuits are genuine contenders for the best approach to sub-threshold FTQC will depend on additional effects, such as leakage errors, not considered in this paper. Such investigation would be interesting future work.

### v2

@@ IV Recovering time-like distance by squeezing two QEC rounds into one
In this section, we show that QEC performance can be improved without using unconditional resets by effectively squeezing two QEC rounds into one. This method, however, requires $\approx 50\%$ additional qubits. We appoint two auxiliary qubits for each stabiliser, as shown in Figure 12(a). Note that the auxiliary qubits of the $X$-/$Z$-type stabilisers are arranged horizontally/vertically so that we avoid so-called “bad hook” errors. Therefore, the hardware needs the Cairo pentagonal connectivity [36]. The main idea is that while we measure the auxiliary qubits of one stabiliser type, we can execute unitary gates on the qubits of the other type and also have plenty of time to obtain two independent measurement outcomes for each stabiliser. Figure 12(b)–(c) illustrates this. For instance, the unitary parts of this circuit for a $Z$-type stabiliser are as follows. We use Roman numerals for qubit labels and numerical labels for protocol steps. First, apply two layers of CZ gates: step (1) $CZ_{i,iii}$ and $CZ_{ii,iv}$; then step (2) $CZ_{i,v}$ and $CZ_{ii,vi}$. Then, swap the auxiliary qubits, so step (3) $SWAP_{i,ii}$. Finally, repeat the two layers of CZ gates: step (5) $CZ_{i,iii}$ and $CZ_{ii,iv}$; then step (6) $CZ_{i,v}$ and $CZ_{ii,vi}$. In this way, we entangle the two auxiliary qubits with the data qubits independently and so measuring each provides independent stabiliser measurements. Since executing these unitary gates takes only $400$ ns in total (when compiled to the gates of Table 1), one QEC round with this circuit takes $1000$ ns. This is only $160$ ns longer than for the standard no-reset syndrome extraction; however, we obtain each stabiliser outcome twice.
@@
We simulated these circuits in order to compare the performance to the previously discussed syndrome extraction circuits. We applied the same principles for the number of shots as in Sections II and III. We present the quantum memory results in Figure 14 and see that again the performance is slightly diminished from the standard no-reset case, though not as much as with the error-spreading circuits. We attribute this to the increased number of two-qubit gates used in these syndrome extraction circuits. However, as we see from the stability time overhead plots in Figure 15(b), we need fewer QEC rounds for the round-squeezing circuit to avoid undetectable time-like logical failures and so we need fewer QEC rounds for quantum memory as well, making the gap in Figure 14 smaller. For instance, with $p=0.001$ we have $R_{\mathrm{sqz}}\approx 0.65$. Since execution times per QEC round differ, we need only $\approx 55\%$ of the number of QEC rounds with the round-squeezing circuit – slightly more than half as many as for the standard no-reset case. Based on this, we can conclude that this circuit improves on the QEC performance of both the standard and error-spreading no-reset circuits, though with the use of additional physical qubits.
From comparing Figure 15(b) and Figure 6(b), we can further conclude that this circuit is a good alternative to implementing unconditional resets, even with instantaneous reset, at least for the range of physical error parameters we consider. We expect that, with much lower physical error probability, this would no longer be the case.
@@
Figure 15: Comparison of stability experiment performances in the no-reset scheme using standard (triangles and dashed lines) and round-squeezing (diamonds and dotted lines) circuits. Colours indicate different physical error probabilities: $p=10^{-2}$ (blue), $p=10^{-2.5}$ (orange), $p=10^{-3}$ (purple). (a) Comparison of logical failure probabilities for a $4\times 4$ stability patch. Error bars as in Figure 6(a) are also shown as shaded areas. Note that the standard $p=10^{-3}$ and round-squeezing $p=10^{-2.5}$ plots overlap. (b) The time overhead $R_{\mathrm{sqz}}$ that the round-squeezing circuit requires when compared to the no-reset standard circuit. The shaded areas correspond to $90\%$ confidence intervals.
@@ V Conclusion
Nevertheless, with current best-in-class devices operating close to the QEC threshold and focused more on memory experiments, our simulations indicated that not resetting qubits can result in better performance. This justifies the approach taken in recent proof-of-principle experiments [6, 8] and offers encouragement for near-term experiments to continue this practice. For those keen to avoid developing unconditional-reset technology, we presented two alternative syndrome extraction circuits that provide additional protection against measurement errors. These alternative circuits are interesting near-term experiments and potential long-term solutions. Whether our alternative circuits are genuine contenders for the best approach to FTQC will depend on additional effects, such as leakage errors, not considered in this paper. Such investigation would be interesting future work.
@@
A summary of our results can be found in Table 2, where we present the best-performing scheme in different regimes based on our circuit-level noise simulations. This table can be used to help decide which scheme to use for FTQC.
Physical error probability Reset speed Suggested scheme to use
Fast No reset
High Slow or No reset
not available
Unconditional reset,
Fast error-spreading or
Medium round-squeezing
Slow or Error-spreading or
not available round-squeezing
Fast Unconditional reset
Low Slow or Error-spreading or
not available round-squeezing
Table 2: Summary of the best-performing scheme(s) in different regimes based on our circuit-level noise simulations. This can be used to help decide which scheme to use for FTQC. For the ‘Physical error probability’ column, ‘high’ corresponds to $p\approx 10^{-2}$, ‘medium’ to $p\approx 10^{-2.5}$ and ‘low’ to $p\lessapprox 10^{-3}$. In the ‘Reset speed’ column, ‘fast’ corresponds to $\approx 100$ ns and ‘slow’ to $\approx 500$ ns.